OpenStudy (anonymous):

What is the exact value of sqrt 38 over 2 sqrt 6 ?

OpenStudy (anonymous):

@jiteshmeghwal9

OpenStudy (anonymous):

can you help please

OpenStudy (jiteshmeghwal9):

rationalize it :)

OpenStudy (anonymous):

how

OpenStudy (jiteshmeghwal9):

\[\Huge{\frac{\sqrt{38} \times 2\sqrt{6}}{2\sqrt{6}\times 2\sqrt{6}}}\]

OpenStudy (anonymous):

oh

OpenStudy (jiteshmeghwal9):

so can u do it from here ?

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

sqrt 57 over 6

OpenStudy (anonymous):

im i right ?

OpenStudy (anonymous):

yes you are ^^ good one!

OpenStudy (anonymous):

You can do it in more simpler way..

OpenStudy (anonymous):

how

OpenStudy (anonymous):

\[\frac{\sqrt{38}}{2 \sqrt{6}} \implies \frac{1}{2} \sqrt{\frac{38}{6}} \implies \frac{\sqrt{19}}{2 \sqrt{3}} \times \frac{\sqrt{3}}{\sqrt{3}} \implies \color{red}{\frac{\sqrt{57}}{6}}\]

OpenStudy (anonymous):

i said that already

OpenStudy (anonymous):

You can use: \[\frac{\sqrt{a}}{\sqrt{b}} \implies \sqrt{\frac{a}{b}}\]